OpenStudy (anonymous):

Rational Numbers and Expressions! Which Property is the following 15 + (k + 4) = 15 + (4 + k) Is it a... Closure property of addition? Closure property of multiplication? Commutative property of addition? Commutative property of multiplication? Associative property of addition? Associative property of multiplication? Identity property of addition? Identity property of multiplication? Inverse property of addition? Inverse property of multiplication? or a Distributive property or is the statement false? (I need help. I have 12 questions and I'm having trouble)

Mehek (mehek14):

Associative property of addition

Mehek (mehek14):

because it doesn't matter where you put the parenthesis

OpenStudy (anonymous):

thanks! could you help me with some more?

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

yea sure

OpenStudy (anonymous):

15 + (4 + k) = (15 + 4) +k

Mehek (mehek14):

that's associative property of addition as well

OpenStudy (anonymous):

Thanks! 15 - (4 - k) = (15 - 4) - k

Mehek (mehek14):

choices?

OpenStudy (anonymous):

it's the same choices

Mehek (mehek14):

i don't see anything for subtraction

Mehek (mehek14):

they're all addition or multiplication

OpenStudy (anonymous):

Do you think it might be Distributive property?

Mehek (mehek14):

distributive property = 5(6+3) = 30 + 15

Mehek (mehek14):

so no

OpenStudy (anonymous):

Okay I'll just mark it for reveiw. Could you still help me with the others?

Mehek (mehek14):

yea

OpenStudy (anonymous):

\[\left[ -6 * \left( -\frac{ 1 }{ 6 } \right) \right] n =1 * n\]

Mehek (mehek14):

choices?

Mehek (mehek14):

if it's true or false, it's true

OpenStudy (anonymous):

I think it might be inverse property of multiplication

OpenStudy (anonymous):

what do you think?

Mehek (mehek14):

inverse property of multiplication = \(\dfrac{7}{4}*\dfrac{4}{7}=1\)

OpenStudy (anonymous):

So no?

Mehek (mehek14):

probably

Mehek (mehek14):

for the ones ur not sure, make a separate post and other people will try to help

OpenStudy (anonymous):

Okay. Thanks for your help!

Mehek (mehek14):

yw :)